The quotient of -36 and 9, subtracted from 4 times a number, is -8. What is the number?
Answer by gwendolyn(128)

This word equation translates into:
4x - (-36)/9 = -8
To simplify this, let's first perform the division:
4x - (-4) = -8
Then, we add -4 to both sides of the equation:
4x = -8 + (-4)
4x = -12
Finally, all we have to do is divide both sides by 4 to isolate the variable:
4x/4 = (-12)/4
x = -3
So, we find that the number is -3.